# ObjectFinder: Open-Vocabulary Assistive System for Interactive Object Search by Blind People

### v1

@@ 2.1. Description-based and Detection-based Systems for Blind People
This image displays the components and connectivity of a wearable system. The system consists of smart glasses equipped with a stereo camera, a microphone on the collar, and two buttons on the waist bag for user inputs. The smart glasses also include bone conduction earphones for output. The image further details the internal components of the glasses which are magnified to show the arrangement from right to left: the right earphone, right infrared camera, color camera, left infrared camera, image ASIC, and left earphone. The NVIDIA Jetson Nano, housed in the waist bag, processes the data collected by these components.
2.2. Multimodal Large Language Models for Assistive Technology for Blind People
Be My Eyes (Eyes, 2024) is the pioneering application to collaborate with OpenAI, integrating GPT-4 (Achiam et al., 2023) to enable blind individuals to comprehend the real world and interpret images. VIALM (Zhao et al., 2024) evaluates the existing Large Language Models (LLMs) for navigation instructions, identifying GPT-4 as the state-of-the-art based on both automatic and human evaluation metrics. VIAssist (Yang et al., 2024) identifies six challenges in capturing desirable images for daily requests, such as incomplete target object placement, and offers reshoot suggestions for low-quality images. VQAsk (De Marsico et al., 2024) is an Android application that enables interaction with one’s photographs or the surrounding environment through speech and haptic feedback. VisionGPT (Wang et al., 2024) leverages LLM to identify existing objects in different scenarios and assess the risk. nanoLLaVA (Face, nd) is the existing smallest LLM developed for the deployment on edge devices, e.g. assistive technology for blind people. Waisberg et al. (Waisberg et al., 2024) proposed embedding LLMs in Ray-Ban glasses from Meta to capture the egocentric view of blind people. To enable blind people to access visual arts, Zhang et al. (Zhang, 2024) designed an interface that enables blind individuals to access visual information from images using LLMs and VLMs. GenAssist (Huh et al., 2023) makes image generation accessible by utilizing a sequence of prompts for LLMs. In this work, we develop our ObjectFinder by using GPT-4 to generate the description and answer questions.
In order to generate proper information from LLM, OpenAI offers an extensive tutorial (OpenAI, 2023) on strategies for crafting effective prompts to achieve optimal results. Numerous studies have explored various aspects of prompt engineering. In this context, we present specific prompt engineering projects designed for the blind community. VIALM (Zhao et al., 2024) identifies the importance of prompting LLMs to generate detailed information regarding obstacles and tactile features in navigation instructions. Gadiraju et al. (Gadiraju et al., 2023) investigate how LLMs can perpetuate subtle yet harmful biases towards communities with special needs, identify specific categories of these biases, and propose training models on diverse resources that positively represent individuals with special needs to mitigate such harms. Hao et al. (Hao et al., 2024) customize prompts for blind people for various purposes and eliminate any potential offense. In this work, we crafted prompts using these strategies to enhance interactive interfaces for blind people, improving digital accessibility and inclusivity.
2.3. Interaction Features
Interaction features are crucial for enhancing the usability and effectiveness of assistive technology for blind persons. Here we discuss the in- and output modalities separately. Speech command input provides a more intuitive and human-like interaction without physical touch, exemplified by voice-controlled smart home devices that offer safety and convenience for blind people (Lee et al., 2022b). Voice assistants such as Apple’s Siri and Amazon’s Alexa are widely utilized by the blind community and are deemed beneficial (Oumard et al., 2022b). Moreover, advanced navigation systems employ voice commands to define targets and utilize GPS for localization (Hersh, 2022; Ramadhan, 2018). He et al. (He et al., 2023) enhance the speech command with bone-conducted vibration on head-mounted wearables. However, the efficacy of speech understanding systems is contingent upon high accuracy, which can be significantly impacted by environmental noise and the systems’ limited ability to comprehend diverse vocabularies (Oumard et al., 2022a). To address these issues, some solutions incorporate simple button-based inputs as alternatives to voice commands (Awad et al., 2018; Tahoun et al., 2020; Hersh, 2022). Kamikubo et al. (Kamikubo et al., 2024) found that in their study, most users referred to active methods of using buttons, and few users prefer to use voice commands to ask specific questions. PathFinder (Kuribayashi et al., 2023) is an automated navigation system with four buttons on the handle interface for functionality selection. In our work, we utilize brief voice commands to enable flexible definition of target objects, while employing buttons to streamline the selection of functions.
In the realm of assistive technologies, methods for conveying information serve distinct purposes. Haptic cues, noted for their intuitive reliability, prove effective across indoor and outdoor settings (Bouteraa, 2021). Vibration is widely employed for alerts in wearable devices (Mahmud et al., 2014; Babu D et al., 2021; Bouteraa, 2021), while tactile cues (Pielot et al., 2012; Liu et al., 2021b) and kinetic force (Amemiya and Sugiyama, 2008, 2009; Ando et al., 2012) are utilized for indicating navigation directions with handheld devices. Yang et al. (Yang et al., 2021) explore light cues on hat brims to guide individuals who have light perception. Auditory signals, such as beeps of varying frequencies, play crucial roles in warning systems for users and pedestrians (Guerrón et al., 2020; Kayukawa et al., 2019). Earcons, which provide contextual information about navigated interfaces, enhance user understanding (Perkins School for the Blind, 2023; OrCam Technologies, 2022). Speech remains the predominant output feature on mobile phones (Leporini et al., 2012; Rodrigues et al., 2015) and wearable systems (Bouteraa, 2021; Chen et al., 2021; Boldu et al., 2020; Liu et al., 2021a; Li et al., 2023), offering comprehensive information delivery. Notably, bone-conduction devices are preferred for their comfort and wearability, enabling the hearing of system instructions alongside environmental sounds (Asakura, 2021; Liu et al., 2021a). Our approach integrates earcons and speech output via bone-conduction headphones within a pair of smart glasses to enhance information accessibility.
In the remainder of this paper, we outline our three-stage research process for developing our object search prototype: (1) a user-in-the-loop approach for designing and refining the prototype, (2) conducting interviews with additional participants to understand their target objects and challenges related to object search, and (3) performing exploratory user studies in both public and private settings to compare the prototype with existing description-based and detection-based systems.
3. Phase I: User-in-the-loop Design
In this phase, we engaged in one-on-one co-design with a blind person (P0, see Table 1) to define functions of the first prototype of the ObjectFinder system, and to develop the design of a study with potential users of the system, rooting initial design decision in the lived experience of a blind person. In this phase, we focus on the challenge of combining description-based and detection-based systems as a foundation to address RQ1. Here, we provide an overview of our design approach, detail our prototype development process, and present the resulting system.
Figure 4. Left: The initial flowchart based on use case definition. The text inside the quotation marks (e.g., “Find”) represents voice input keywords to trigger target identification, target confirmation, and function selection. Right: The system architecture includes the interaction steps and multiple functions corresponding to the flowchart. Given a task (e.g., “finding a coffee table in an unknown space”), a target with location and orientation information can be found and output. The visualization is derived from a real example in our user study.
@@ 3. Phase I: User-in-the-loop Design
This figure illustrates the initial flowchart based on the use case definition on the left side and the system architecture on the right side. The flowchart has the following elements: 1. Vocal input: ’Find ¡target¿’ (block circled 1) a. Flows to 2. 2. Confirm ¡target¿? a. If yes, flows to 3. b. If no, flows back to 1. 3. Scan the environment, while continuously judging if the target is found. (block circled 2) a. If yes, flows to 4. 4. Localization information: ¡target¿ ¡distance¿ ¡direction¿ (block circled 3) a. Flows to 5. 5. Voice command to choose route planning, scene description, or open questions. (block circled 4) a. If the voice command ’continue searching’ is given, flows back to 1. On the right side of the figure, the network architecture, including the inputs and outputs of YOLO World and MobileSAM, is depicted. These components are illustrated to demonstrate their functionality in efficient object detection and enhancing the precision of localization information.
3.1. Definition of the Use Case
In a first step, we started with the use case definition, following the principle in (Cockburn, 2000). For this purpose, we organized a workshop that included the blind co-designer (P0), one developer, and two accessibility and usability experts (one of them is a blind person). The use case search for an object, here a socket, in an unfamiliar environment was identified in the workshop. Then, the blind co-designer had the task imagining how he would use a system to look for a socket in an unknown office, while the other participants asked questions that helped the co-designer to define the requirements and the different steps of the search. We formulated the use case in a diagram in Figure 2 for a clear illustration. The co-designer’s expectation suggested decomposition of the use case into $5$ functions: Object Detection (F1), Localization (F2), Route Planning (F3), Scene Description (F4), and Open Questions (F5). Leveraging these categories, we first restructured the interaction features and functions into an initial flowchart, as shown in Figure 4. In the following, we explain assembly of hardware components of ObjectFinder, and we outline implementation of basic functionality into the first prototype.
3.2. Iteration 1: Initial Prototype Design
In this section, we describe the implementation of the initial prototype based on this use case, detailing the hardware and interaction features and the functionality implementation.
3.2.1. Hardware and Interaction Features
Figure 3 presents the system diagram, which comprises a pair of KRVision (KR Vision, [n. d.]) smart glasses coupled with a waist bag. The smart glasses are outfitted with a RealSense R200 RGB-Depth camera, facilitating real-time RGB and depth frame acquisition in an egocentric perspective. Additionally, a bone conduction headset is incorporated, enabling auditory output while maintaining perception of environmental sounds. In the waist bag, a Jetson Nano, a compact and powerful processor, is utilized for efficient data processing, accompanied by a power bank for energy supplementation. The waist bag features two buttons which can be programmed and used for volume adjustment initially. A microphone is attached to the collar for audio input. As shown in Figure 4, participants simply say commands with keywords to trigger target identification, target confirmation, and function selection.
Figure 5. Rectified flowchart. Left: The optimized version of the flowchart for the open-vocabulary interactive ObjectFinder. Right: Steps added to the ObjectFinder baseline. The object detection model in the baseline can only identify a list of objects, the added steps assist in locating the other objects through egocentric scene descriptions.
@@ 3.2.1. Hardware and Interaction Features
On the left side of the figure is the rectified flowchart, and on the right side are the different steps between ObjectFinder and ObjectFinder_Base. The rectified flowchart includes the following steps: 1. Vocal input: ’Find ¡target¿’ (block circled 1) a. Flows to 2. 2. Confirm ¡target¿? a. If yes, flows to 3. b. If no, flows back to 1. 3. Scan the environment while continuously judging if the target is found. (block circled 2) a. If yes, flows to 4. b. If time out, flows to 5. 4. Localization information: ¡target¿ ¡distance¿ ¡direction¿ (block circled 3) a. Flows to 6. 5. Press 1 and 2 to trigger continued searching and scene description. (block circled 5) a. If 1 is pressed, flows back to 1. 6. Press 1, 2, 3, and 4 to trigger route planning, scene description, continued searching, and open questions. (block circled 4) a. If 1 is pressed, flows back to 1. b. Otherwise, conduct the function and flow to 7. 7. Press 1, 2, and 3 to trigger continued searching, repeat, and either scene description or open questions. (block circled 6) a. If 1 is pressed, flows back to 1. b. Otherwise, remain in this block. On the right, the different steps between blocks circled 1 and 2 are illustrated: 1. Confirm ¡target¿? a. If yes, flows to 2. 2. Judge if the target is in COCO2017. a. If yes, flows to circled 2. b. If no, flows to 3. 3. Press 1 to continue searching, press 2 for scene description.
3.2.2. Functionality Implementation
To achieve the aforementioned open-vocabulary interaction, all five functions are primarily realized by a combination of models: YOLO-World (Cheng et al., 2024) and MobileSAM (Zhang et al., 2023) as Open-Vocabulary Object Detection (OVOD), GPT-4 (Achiam et al., 2023) as Multimodal Large Language Model (MLLM). As the initial flowchart and architecture of the system are shown in Figure 4, we detail the implementation of the individual function in the following.
Object Detection (F1).
The recently proposed YOLO-World (Cheng et al., 2024), employing a prompt-then-detect strategy, allows users to define their customized vocabularies for the object detector. In our prototype, users specify the target object via a voice command “Find <target>”. The prototype then verifies whether the object exists within the predefined vocabulary list. If absent, the new target word is embedded and saved in the model’s vocabulary for future reference, a process requiring approximately $20$ seconds for initialization. To enhance efficiency, we have pre-populated the vocabulary with a list of common everyday objects. We utilize the large version of YOLO-World, which performs inference in $0.5$ seconds on our system. The key frame where the object is detected is retained for subsequent analysis.
Localization (F2).
@@ 3.2.2. Functionality Implementation
Our study utilized an Object-Distance-Direction format for localization as proposed by Constantinescu et al. (Constantinescu et al., 2022). Directions are presented in a clockwise orientation and distances in meters. Initial distances were estimated using YOLO-World’s average bounding box distances due to the pending release of its segmentation version. These bounding boxes, which need to overlap ground truth by at least $50\%$, might inaccurately include significant background areas in calculations (Liu et al., 2023). To enhance accuracy, MobileSAM (Zhang et al., 2023), the compact version of SAM (Kirillov et al., 2023), was later added to generate segmentation masks using the bounding boxes as prompts post YOLO-World (Cheng et al., 2024).
Multimodal Large Language Model Functions (F3-F5).
@@
According to VIALM (Zhao et al., 2024), GPT-4 is the state-of-the-art for guiding blind people, excelling in both human (correctness, actionability, fluency) and automated evaluations. Thus, we chose GPT-4 for implementing MLLM functions (F3-F5). Effective prompt engineering is crucial for enhancing large language models’ utility and accessibility for blind users. Our system prompt is concrete, incorporating role, tone, and response length, guided by OpenAI’s Prompt Engineering tutorial (OpenAI, 2023). Response length, aligned with social media alt-text limits (100-500 characters) (for the Blind, 2024), is set to the maximum to cater to user preferences for vivid responses. For navigation instructions, Wang et al. (Wang et al., 2023) suggest including position, count, size, color, material, and shape. We developed two user prompts: one for route planning (F3) and another for scene description (F4), with voice input for open questions (F5). The concrete prompts are provided in the supplementary material.
3.3. Iteration 2: System Optimization and Refinement
We conducted four rounds of preliminary half-hour user exploration, inviting P0 to explore the prototype freely in a demo room that integrates two scenarios: a private setting (i.e., a living room) and a public setting (i.e., an office). The layouts of these scenarios are shown in Figure 7(b). P0 could search for anything and trigger any functions he wanted. The wearable prototype is designed to free the hands of blind people to hold a white cane. However, it was difficult to determine which information was acquired from our prototype in the small scenarios. Therefore, we decided that participants in the user study will not hold a white cane while exploring the system. An assistant who is not related to this project should help to ensure the safety of the participants. The problems P0 encountered and the improvements of the prototype are summarized in two aspects: interaction features and flowchart. The resulting flowchart is formulated in Figure 5.
3.3.1. Interaction Features
An acoustic signal recognition module triggers steps like target definition, target confirmation, and option selection in our initial flowchart. Due to language redundancy, P0’s voice input wasn’t always fully recognized or understood. Inspired by call forwarding, we adapted the system to replace most voice commands with button selections, except for target and open question definitions. Consequently, the two buttons on the waist bag, labeled “1” and “2”, handle confirmations and selections: “1” for “yes”, “2” for “no”, and combinations of “1”, “2”, “12”, or “21” for choosing among up to four options. We omitted double-clicking due to the buttons’ limited key travel. We employed a continuous low-frequency beep to indicate the that system is in operation and enhance user’s sense of safety. Additionally, taking cues from Orcam MyEye2 (OrCam Technologies, 2022), we integrated earcons for target detection and input readiness, with beeps at $1Hz$ for system activity and $3Hz$ for model loading and MLLM responses. After confirming targets, users should move their heads to scan and stand still following the earcons. A misalignment issue between the user’s head and body prompted us to add, “Please align your body with your head’s direction.” to each MLLM output to clarify the expected posture.
@@ 3.3.1. Interaction Features
P0 Male 30-39 Light perception, since about 2004 Be My Eyes, Seeing AI
P1 Female 20-29 Light perception, since about 2022 Be My Eyes, Seeing AI
P2 Male 50-59 Fully blind, since birth Be My Eyes, Seeing AI
P3 Male 20-29 Fully blind, since birth Be My Eyes
@@
Table 1. Participant demographic statistics in Phase II interview with target group.
3.3.2. Flowchart Refinement
Participant P0 noted that route planning (F3) was a unique MLLM feature distinguishing our ObjectFinder from Be My Eyes (Eyes, 2024) and frequently activated this feature. However, remembering the instructions after a single playback proved challenging, so the instructions needed to be repeated. Additionally, P0 might lose orientation after taking several steps toward the target, necessitating additional scene descriptions. To address potential confusion regarding the route, we encourage users to search for the target again. Consequently, we provided options to continue searching, repeat, and scene descriptions in block 6, following the route planning section in block 4. We also introduced open questions in block 6 after scene description (F4) and open question (F5) in block 4 to solicit more detailed feedback. In scenarios where the target object was absent, we added block 5 to inform users of available objects. (see Figure 5)
4. Phase II: Interview Study
During Phase II, we conducted interviews with $8$ participants to understand the technology they use and to identify challenges they face when searching for objects.
4.1. Method
To better understand the target objects, we asked participants about the items they most frequently or laboriously search for in their daily lives, and which objects they would be interested in locating using the system, such as furniture or smaller items. To gauge whether they require inspiration for unexpected targets (C2), we inquired about the objects they would expect to find in an unfamiliar office or living room, and what items they might not recognize unless informed of their existence. We also explored the advantages and disadvantages of the technologies and methods participants currently use for object search or scene understanding. Additionally, we specifically asked whether they could match photos taken by a description-based system, such as Be My Eyes, with the actual environment.
4.1.1. Participants and Procedure

### v2

@@ 3.1. Design Goals
This image displays the components and connectivity of a wearable system. The system consists of smart glasses equipped with a stereo camera, a microphone on the collar, and two buttons on the waist bag for user inputs. The smart glasses also include bone conduction earphones for output. The image further details the internal components of the glasses, which are magnified to show the arrangement from right to left: the right earphone, right infrared camera, color camera, left infrared camera, image ASIC, and left earphone. The NVIDIA Jetson Nano, housed in the waist bag, processes the data collected by these components.
3.2. Envisioned Scenario
In an initial step, we began by defining the use case according to the principles outlined in (Cockburn, 2000). To achieve this, we conducted a workshop involving the blind co-designer, a developer, and two experts in accessibility and usability, one of whom is blind.
The co-designer presented a use case for object search: searching for a socket in an unfamiliar office. We refined the use case regarding the interaction sequence between the co-designer and the wearable object search system as the envisioned scenario (Carrol, 1995), serves as the basis for our system design, as illustrated in Figure 2 and depicted as follows:
Martin enters an unfamiliar office, his phone battery depleted. In need of power, Martin activates an object search system and commands it to “Find a socket”. The system acknowledges the command, and ensures Martin that it has understood the request. After Martin confirms, the system signals with a sound, indicating readiness to begin searching.
As Martin scans the room through the system, he prefers not to be bombarded with information about every detected object; instead, he wants the system to announce only when it detects a socket. Upon identifying a socket, the system provides feedback on its egocentric location, including distance and direction, as well as its allocentric relationship with points of interest.
Using this information, Martin evaluates the suitability of the socket. The first socket detected, located near a trash bin about 4 meters away at a 10 o’clock direction, is deemed inconvenient because Martin intends to work at a workstation. Therefore, he continues his search for a more suitably placed socket.
Eventually, a socket near a workstation, just 3 meters away in the 2 o’clock direction, catches Martin’s interest. He requests more details about this socket, such as directions to reach it and its height on the wall. The system advises Martin to navigate around obstacles, guiding him with instructions like, “Walk around the clutter…”
Utilizing his cane to detect and avoid clutter, Martin reaches the workstation situated to his front right and successfully charges his phone using the nearby socket.
3.3. Hardware and Interaction Features
According to the related work (Herskovitz et al., 2023; Gamage et al., 2023) and the envisioned scenario with the co-designer, a pair of glasses is assumed to be preferred over a smartphone for an object search system. We utilize the following hardware to implement this system. Figure 3 presents the system diagram, which comprises a pair of KRVision smart glasses (KR Vision, [n. d.]) coupled with a waist bag. The smart glasses are outfitted with a RealSense R200 RGB-Depth camera, facilitating real-time RGB and depth frame acquisition in an egocentric perspective. Additionally, a bone conduction headset is incorporated, enabling auditory output while maintaining perception of environmental sounds. In the waist bag, an NVIDIA Jetson Nano, a compact and powerful processor, is utilized for efficient data processing, accompanied by a power bank for energy supplementation. The waist bag features two buttons that are programmed for target confirmation and function selection. A microphone is attached to the collar for audio input: participants simply speak commands to specify targets or ask questions after triggering open questions. The initial version of ObjectFinder is only for the prototype. We aim to further integrate all hardware components more compactly to improve the user’s daily experience in real-world use, e.g., with Ray-Ban glasses (Waisberg et al., 2024).
3.4. Function Implementation
According to the scenario envisioned in Section 3.2, we have decomposed the object search into five functions: Object Detection (F1), Localization (F2), Route Planning (F3), Scene Description (F4), and Open Questions (F5). To integrate these five functions, we define three modules in the pipeline: user specifies the target, system detects the target, and system generates feedback, see Figure 4.
3.4.1. Module 1: User specifies target.
When the system is turned on, the frame of the scenario is captured by the smart glasses automatically. GPT-4 then generates a list of objects based on the frame to initialize YOLO-World. The user specifies the target object using the command frame, “Find ¡target¿”. After receiving the command, the system will repeat the target object: “You want to find ¡target¿, please confirm.”. For confirmation, the user should press the button with a sticker on the waist bag, while the other button is for respecification. Speech-to-text is processed by Google Speech Recognition API, and text-to-speech is handled by the pyttsx3.
The relationships between the specified target objects and the object classes in YOLO-World are categorized into three types: match, related to, and unrelated to, as shown in Figure 5. We define match as instances where the string of one item in the object list appears within the target object. This is important because recorded speech may sometimes be unclear due to user rephrasing, such as “Find the chair, no, office chair.” and the surrounding conversations.
For other cases, the specified target object and each object class in YOLO-World are tokenized and embedded with all-MiniLM-L6-v2 (Reimers and Gurevych, 2021). The cosine similarity between the embeddings of the target object and the object classes of YOLO-World is calculated. If the similarity score between the target object and any object class in YOLO-World is at least 0.8, a threshold we have set based on empirical data, the target object is deemed to be related to the object class with which it shares the highest similarity in the current list of YOLO-World objects. If the target object is deemed related to an existing item in YOLO-World, this item will be used for subsequent object detection searches, thereby eliminating the need to reinitialize the system. Conversely, if no similarity between the objects in the list and the target object exceeds the threshold, then the target object is considered unrelated to the current object list of YOLO-World. In this case, YOLO-World should be reinitialized with the object list updated to include the new target object. During the YOLO-World initialization process, a $3$ Hz beep is played in the background to reassure the user that the system is still operating.
Figure 4. ObjectFinder system architecture integrates five functions into three modules for interactive object search. (a) Initially, an open-vocabulary object detector, e.g. YOLO-World, is initialized with a list of objects extracted from a scenario capture, allowing the user to identify a target object. If the target is not on the list, the object detector is reinitialized. (b) The user scans the environment. If the target is detected, localization information is provided in real-time. If not, the user can trigger scene understanding to identify what exists in the scenario. (c) The user may activate a sub-branch to obtain further information based on their intent using a multimodal large language model. (d) If the user discovers other objects of interest or becomes disoriented, they can reorient themselves to locate the target.
@@ 3.4. Function Implementation
The image shows a system designed to help blind users search for objects. It begins with the user capturing a scene and specifying a target (section a). The system initializes object detection using YOLO-World and compares it with the target. In the perception module (section b), detected objects are localized, or the system moves to scene understanding if detection times out. Based on intent (section c), the system supports navigation (e.g., route planning) or perception tasks (e.g., answering questions or refining the search). Feedback loops allow continuous user-system interaction for effective object search.
3.4.2. Module 2: System detects target.
Figure 5. Initialization with target identification: (a) The list of detectable objects in YOLO-World is initialized with the first capture of the scenario. The target objects can be categorized into three types: (b) match, where the object matches an item in the list; (c) related, where the object is related to one item in the list (e.g., “couch” is related to “sofa” with 0.85 similiarity); and (d) unrelated, where the object does not relate to any item in the list. In cases where the object is unrelated, the list is updated by adding the target to it.
@@
The figure illustrates how a similarity-based matching process works for object search by blind users. On the left, a room scene includes a sofa, cushions, lamp, paintings, and a fan. GPT-4 generates an object list from this scene. Three example target queries are shown: ”fan,” ”couch,” and ”teapot.” The target ”fan” directly matches the object list, requiring no similarity calculation. For ”couch,” the system computes similarities with existing objects: sofa (0.85), cushions (0.61), lamp (0.36), paintings (0.20), and fan (0.27). Since ”sofa” has the highest similarity and exceeds the threshold of 0.8, it is selected without needing to reinitialize YOLO-World. In contrast, ”teapot” has low similarity scores across all objects—sofa (0.25), cushions (0.25), lamp (0.33), paintings (0.21), and fan (0.19)—none exceeding the threshold. As a result, the system triggers YOLO-World reinitialization using the original object list to find the teapot.
After YOLO-World is initialized with the target object, the user will hear an earcon to signal the start of scanning. The system successfully detects the target (F1) when the confidence level of its bounding box exceeds the empirically set threshold of $0.3$. The system captures a key frame that includes both RGB and depth information. At the same time, another earcon sounds, signaling the user to pause and orient themselves toward the target object. This frame is used to calculate localization information (F2) and to query for further intent-based, long-text feedback. If the target object is not detected within a time limit of $45$ seconds, it is considered absent in the scenario. The user can choose to activate scene understanding (F4) or re-specify the target object.
@@
If the target is detected successfully, the egocentric localization information will be calculated using the keyframe and delivered in the format Object-Distance-Direction, as proposed by Constantinescu et al. (Constantinescu et al., 2022), as illustrated in Figure 6. Egocentric information is presented in a clockwise orientation and distances in meters. Calculating distance using a bounding box is inaccurate. The bounding box for object detection is considered accurate if it overlaps with the ground truth by at least $50\%$. However, this criterion might result in the bounding box inaccurately encompassing significant background areas (Liu et al., 2023). To enhance accuracy, MobileSAM (Zhang et al., 2023), the compact version of SAM (Kirillov et al., 2023), was later added to generate segmentation masks $M$ using the bounding boxes as prompts post YOLO-World (Cheng et al., 2024). The distance of the object is calculated as the average depth from the key frame’s depth map, masked by $M$.
(1) $Distance=\frac{\sum(\text{depth\_map}\odot M)}{\sum M}$
To determine the clockwise direction, we use the center of the bottom edge of the frame, denoted as $(x_{c},y_{c})$, as the clock’s center. The angle $\theta$ between the center of the bounding box $(x_{\text{bbox}},y_{\text{bbox}})$ and the clock’s center, relative to the bottom edge, is calculated as follows:
(2) $\theta=\arctan\left(\frac{y_{c}-y_{bbox}}{x_{c}-x_{bbox}}\right)\times\frac{180}{\pi}$
Then, $\theta$ in the range $(-90^{\circ},90^{\circ})$ is mapped to the clock positions from 9 to 3 o’clock.
3.4.3. Module 3: System generates feedback.
When generating feedback, the user is halted by an earcon and oriented towards the target object. Simultaneously, the keyframe capturing the user’s egocentric view, which includes the target object, is sent to the MLLM to produce long-text feedback.
If the user considers the detected candidate to potentially be the target object, they may wish to learn more based on their intent. During the refinement iteration with the co-designer, we observed that he primarily had two intents regarding detected objects: navigating to functional objects for interaction, such as finding a charger to charge a smartphone, and perceiving regions of interest, such as identifying items on a tabletop without interacting with them. This was followed by a request for further details. Therefore, we have currently implemented branches for these intentions in this module, with the possibility of adding more in the future.
Generating feedback based on user intent.
In the process of generating intent-based feedback, the system initially uses the keyframe for the first query, while subsequent queries are based on the current egocentric frame of the user. In the navigation branch, the user initiates route planning (F3). After moving several steps, the target object may no longer be in the field of view, which could prevent further route planning. Consequently, the user has the option to either repeat the instruction or trigger a scene description (F4) for orientation. For example, if the user is aware that there is a desk on the route to the fan and understands from the scene description (F4) that the desk is directly in front of them, they will know “I’m getting close to the fan.” If the user still feels lost, they can revert to the target specification to relocate the target object. In the perception branch, the user can opt to use scene description (F4) to detail the surroundings of the target object, or directly ask open questions (F5) to engage in a conversation about the target object and its surroundings over several rounds. As the co-designer always discovered objects of interest or rejected the candidate after learning about the detailed information surrounding it, the user can respecify the target object at any step with ObjectFinder.
Optimizing interaction features.
Following the approaches of (Awad et al., 2018; Tahoun et al., 2020; Hersh, 2022), we have programmed the two buttons on the waist bag to select functions. We opted for this method over speech commands. Despite the flexibility, it is susceptible to environmental noise and the system’s comprehension limitations (Oumard et al., 2022). Speech input is used only for target specification and dialogue in open questions (F5) for efficient object search.
Figure 6. Object detection and localization: Each video frame is processed by YOLO-World to detect key frames in which the confidence level of the bounding box around the target object exceeds a certain threshold. Subsequently, the segmentation map generated by the bounding box, combined with the depth map, is used to provide precise localization information, including distance and direction.
[…]
The figure illustrates a pipeline for object detection and localization. A video stream is processed frame by frame using YOLO-World, which identifies objects and highlights keyframes where the target object is detected with high confidence (e.g., a coffee table with a confidence score of 0.88). The selected keyframe, including the bounding box, is passed into MobileSAM, a segmentation module. Within MobileSAM, the image and bounding box are processed through an image encoder and prompt encoder, respectively, then combined in a mask decoder to produce a segmentation map highlighting the target object. This segmented output is then fused with a depth map generated from the corresponding video frame to extract accurate localization information, such as the object’s distance and direction from the camera.
Enhancing feedback accessibility through prompt engineering.
@@
According to VIALM (Zhao et al., 2024), GPT-4 is the state-of-the-art for guiding blind people, excelling in both human (correctness, actionability, fluency) and automated evaluations. Thus, we chose GPT-4 for implementing MLLM functions (F3-F5). Effective prompt engineering is crucial for enhancing large language models’ utility and accessibility for blind users. Our system prompt is concrete, incorporating role, tone, and response length, guided by OpenAI’s Prompt Engineering tutorial (OpenAI, 2023). The response length, aligned with the alt text limits of social media (100-500 characters) (for the Blind, 2024), is set to the maximum to accommodate user preferences for vivid responses. The co-designer scans the environment by turning his head rather than his body. When a target object is captured, he always pauses in his current posture, with his head and body often misaligned. Therefore, if the system provides egocentric instructions like “Please walk two steps forward.” or “The desk is in front of you,” it can lead to confusion. To resolve this ambiguity, we precede each response from the MLLM with the instruction, “Please align your body with the direction of your head.” This adjustment helps tailor the system’s feedback to accommodate any user’s scanning strategy. For route planning, we specifically consider that instructions based on steps are easier for blind people to understand, as suggested by (Gamage et al., 2023). Additionally, we provide instructions using landmarks rather than turn-by-turn directions, as per (Jain et al., 2023b). The system prompt and two user prompts for route planning (F3) and scene description (F4) are detailed in the supplementary material.
Table 2. Demographics of participants. P0 is the co-designer who helped to adapt the system to the needs of the target group. P1-P8 were participants of the user study.
@@
P0 Male 30-39 Light perception, since about 2004 BeMyAI, Seeing AI
P1 Female 20-29 Light perception, since about 2022 BeMyAI, Seeing AI
P2 Male 50-59 Fully blind, since birth BeMyAI, Seeing AI
P3 Male 20-29 Fully blind, since birth BeMyAI
@@
Figure 7. Simplified user study procedure focusing on two target objects: a large piece of furniture (an office chair) and a smaller object (a plate of cookies). The user walks into an unfamiliar environment, the office in this example. To sit in front of a desk, he or she should find an office chair first and navigate to it. Then the user sits on the office chair and defines the desk as the region of interest. The user will know what is on the desk through the scene description function, a plate of cookies in this example, and get additional information through open questions.